OpenStudy (anonymous):

express in terms of i. 2 and square root of -16

OpenStudy (anonymous):



OpenStudy (anonymous):

\( 2 \sqrt{-16} = 2 \sqrt{16(-1)} = 2 \sqrt{16}\sqrt{-1} = 8i\) :-)